Budgeted Income Statement and Supporting Budgets

The budget director of Gold Medal Athletic Co., with the assistance of the controller, treasurer, production manager, and sales manager, has gathered the following data for use in developing the budgeted income statement for March 2016:

Estimated sales for March:

Batting helmet 1,200 units at $40 per unit
Football helmet 6,500 units at $160 per unit

Estimated inventories at March 1:

Direct materials:
Plastic 90 lbs.
Foam lining 80 lbs.
Finished products:
Batting helmet 40 units at $25 per unit
Football helmet 240 units at $77 per unit

Desired inventories at March 31:

Direct materials:
Plastic 50 lbs.
Foam lining 65 lbs.
Finished products:
Batting helmet 50 units at $25 per unit
Football helmet 220 units at $78 per unit

Direct materials used in production:

In manufacture of batting helmet:
Plastic 1.20 lbs. per unit of product
Foam lining 0.50 lb. per unit of product
In manufacture of football helmet:
Batting helmet 3.50 lbs. per unit of product
Football helmet 1.50 lbs. per unit of product

Anticipated cost of purchases and beginning and ending inventory of direct materials:

Plastic $6.00 per lb.
Foam lining $4.00 per lb.

Direct labor requirements:

Batting helmet:
Molding Department 0.20 hr. at $20 per hr.
Assembly Department 0.50 hr. at $14 per hr.
Football helmet:
Molding Department 0.50 hr. at $20 per hr.
Assembly Department 1.80 hrs. at $14 per hr.

Estimated factory overhead costs for March:

Indirect factory wages $86,000
Depreciation of plant and equipment 12,000
Power and light $4,000
Insurance and property tax 2,300

Estimated operating expenses for March:

Sales salaries expense $184,300
Advertising expense 87,200
Office salaries expense 32,400
Depreciation expense—office equipment 3,800
Telephone expense—selling 5,800
Telephone expense—administrative 1,200
Travel expense—selling 9,000
Office supplies expense 1,100
Miscellaneous administrative expense 1,000

Estimated other income and expense for March:

Interest revenue $940
Interest expense 872

Estimated tax rate: 30%

2. Prepare a production budget for March. Enter all amounts as positive numbers.

Gold Medal Athletic Co.
Production Budget
For the Month Ending March 31, 2016
Units
Batting helmet Football helmet
Expected units to be sold
Plus desired inventory, March 31, 2016
Total
Less estimated inventory, March 1, 2016
Total units to be produced


3. Prepare a direct materials purchases budget for March. Enter all amounts as positive numbers.

Gold Medal Athletic Co.
Direct Materials Purchases Budget
For the Month Ending March 31, 2016
Plastic Foam Lining Total
Units required for production:
Batting helmet
Football helmet
Plus desired units of inventory, March 31, 2016
Total
Less estimated units of inventory, March 1, 2016
Total units to be purchased
Unit price $ $
Total direct materials to be purchased $ $ $


4. Prepare a direct labor cost budget for March. Enter all amounts as positive numbers.

Gold Medal Athletic Co.
Direct Labor Cost Budget
For the Month Ending March 31, 2016
Molding
Department Assembly
Department Total
Hours required for production:
Batting helmet
Football helmet
Total
Hourly rate $ $
Total direct labor cost $ $ $


5. Prepare a factory overhead cost budget for March. Enter all amounts as positive numbers.

Gold Medal Athletic Co.
Factory Overhead Cost Budget
For the Month Ending March 31, 2016
Indirect factory wages $
Depreciation of plant and equipment
Power and light
Insurance and property tax
Total $


6. Prepare a cost of goods sold budget for March. Work in process at the beginning of March is estimated to be $15,300, and work in process at the end of March is desired to be $14,800. Enter all amounts as positive numbers.

Gold Medal Athletic Co.
Cost of Goods Sold Budget
For the Month Ending March 31, 2016


Direct materials:


Cost of direct materials available for use

Cost of direct materials placed in production


Total manufacturing costs
Total work in process during the period

Cost of goods manufactured
Cost of finished goods available for sale

Cost of goods sold


7. Prepare a selling and administrative expenses budget for March. Enter all amounts as positive numbers.

Gold Medal Athletic Co.
Selling and Administrative Expenses Budget
For the Month Ending March 31, 2016
Selling expenses:
Sales salaries expense
Advertising expense
Telephone expense—selling
Travel expense—selling
Total selling expenses
Administrative expenses:
Office salaries expense
Depreciation expense—office equipment
Telephone expense—administrative
Office supplies expense
Miscellaneous administrative expense
Total administrative expenses
Total operating expenses


8. Prepare a budgeted income statement for March. Enter all amounts as positive numbers.

Gold Medal Athletic Co.
Budgeted Income Statement
For the Month Ending March 31, 2016



Operating expenses:


Total operating expenses
Income from operations
Other income:

Other expenses:

Income before income tax

Net income
Answer #1
Concepts and reason

Budget: Budget is a quantitative estimate of the work to be performed. It acts as a tool for the activities to be performed in the future. It is the process of planning before acting.

Sales revenue: The amount earned by selling goods or services is known as sales revenue. This is the main source of revenue for a company. This is usually referred to as gross sales revenue. Gross sales revenue less returns, allowances, and discounts are known as net sales.

Cash sales: Cash received once the activity of sales commenced is known as cash sale. It is payment received with ready cash during the sale. Sellers offer discount for such sale to encourage cash sales among customers or buyers.

Credit sales: Credit sales are the sales done and the income or revenue is accrued but not received immediately. Cash is received during the future period for the sales. The customer is thus known as debtor.

Inventory: Inventory refers to the goods purchased by the company from manufacturers and resell to the customers. Transportation charges at the time of purchase, storage and insurance cost etc. are included in the merchandise inventory account. Inventory is one of the important current assets of the company

Fundamentals

Direct materials: Direct materials are the raw materials which are directly related with the production of the goods

Direct labor cost: It refers to the cost of providing wages to the workers who are directly associated with the production of goods or services rendered to the customers. The cost of direct labour includes the wages, payroll taxes, and all the benefits sponsored by the manufacturer. The format for preparing direct labor cost budget is

Total
G Co
Direct Labor cost Budget
For the Month ending March 31, 2016
Molding Assembly
Hours required for production
Batt

Factory overhead: Factory overhead costs are the indirect cost incurred to manufacture a product. It is also known as work cost. It is classified into variable factory overheads and fixed factory overheads. Variable factory overheads only include the variable cost and fixed factory overhead include the fixed costs. The format for preparing factory overhead budget is

G Co
Factory overhead Budget
For the Month ending March 31, 2016
Indirect factory wages
Depreciation of plant and equipment

Production budget: Production budget is a budget prepared before the production commences. It shows the forecast about all the details related to the production. The format for production budget is

G Co
Production Budget
For the Month ending March 31, 2016
Particulars
Units
Batting helmet Football helmet
Expected units

Material purchase Budget: Material purchase Budget is a budget prepared to estimate the requirement of materials used for production. It is prepared before the purchase of materials and production process

Foam
Total
G Co
Material Purchase Budget
For the Month ending March 31, 2016
Particulars
Plastic
Units required for product

Cost of goods sold: The costs which are incurred by a business to sell products in a particular period are called as cost of goods sold. It is also known as cost of sales. It is considered to be as expense of current period. The format for preparing cost of goods sold budget is

G Co
Cost of Goods sold Budget
For the month ending March 31, 2016
Finished goods inventory, March 1, 2016
Work in process

Selling and administrative expenses budget: It is a budget prepared to calculate the total operating expenses that is concerned with the total of selling and administrative expenses budget. The following is the format for the selling and administrative expenses budget.

G Co
Selling and Administrative expenses Budget
For the month ending March 31, 2016
Particulars
Amount Amount
Selling expen

Net income: This is the income of a company which reports all the revenues earned and expenses incurred by the company over a specific period of time. It is the income after adjusting the profit and losses. It is also known as operating income and earnings before interest and taxes. The format for preparing budgeted income statement is as follows

G Co
Budgeted Income statement
For the month ending March 31, 2016
Revenue from sales
Less: Cost of goods sold
Net Income
L

1.2)

Computation of Production Budget:

G Co
Production Budget
For the Month ending March 31, 2016
Particulars
Units
Batting helmet Football helmet
Expected units

Therefore, units to be produced are 1,210 batting helmets and 6,480 football helmets.

1.3)

Computation of Materials purchases Budget:

G Co
Material Purchase Budget
For the Month ending March 31, 2016
Particulars
| Plastic Foam Lining Total
Units required fo

Therefore, total cost of direct material to be purchased is $185,792.

Workings:

Calculation of units required for production:

Calculation of units of batting helmet and football helmet:

UnitsofBattinghelmet=UnitsofProduction×NoofPlasticmaterials=1,210×1.20=1,452unitsUnitofFootballhelmet=6,480×3.50=22,680units\begin{array}{c}\\{\rm{Units of Batting helmet = Units of Production}} \times {\rm{No of Plastic materials}}\\\\{\rm{ = 1,210}} \times {\rm{1}}{\rm{.20}}\\\\{\rm{ = 1,452 units}}\\\\{\rm{Unit of Football helmet = 6,480}} \times {\rm{3}}{\rm{.50}}\\\\{\rm{ = 22,680 units}}\\\end{array}

Therefore, units of batting helmet are 1,452 units and units of football helmet are 22,680 units.

1.4)

Computation of Direct labor cost Budget:

G Co
Direct Labor cost Budget
For the Month ending March 31, 2016
Molding Assembly Total
Hours required for production
Batt

Therefore, total direct labor cost is $241,406.

Workings:

Calculation of Hours required for production:

Calculation for molding department:

Battinghelmet=Unitsproduced×Requiredhoursperunit=1,210×0.20=242HoursFootballhelmet=6,480×0.50=3,240Hours\begin{array}{c}\\{\rm{Batting helmet = Units produced}} \times {\rm{ Required hours per unit}}\\\\{\rm{ = 1,210}} \times 0.{\rm{20}}\\\\{\rm{ = 242 Hours}}\\\\{\rm{Football helmet = 6,480}} \times 0.50\\\\ = 3,240{\rm{ Hours}}\\\end{array}

Therefore, hours required for molding department is 242 hours for batting helmet and 3,240 hours for football department.

Calculation for assembly department:

Hoursrequiredforproduction=Unitsproduced×RequiredhoursperunitBattinghelmet=1,210×0.50=605hoursFootballhelmet=6,480×1.80=11,664hours\begin{array}{c}\\{\rm{Hours required for production = Units produced}} \times {\rm{ Required hours per unit}}\\\\{\rm{Batting helmet = 1,210}} \times {\rm{0}}{\rm{.50}}\\\\{\rm{ = 605 hours}}\\\\{\rm{Football helmet = 6,480}} \times {\rm{1}}{\rm{.80}}\\\\{\rm{ = 11,664 hours}}\\\end{array}

Therefore, hours required for assembly department is 605 hours for batting helmet and 11,664 hours for football department.

1.5)

Computation of Factory overhead Budget:

G Co
Factory overhead Budget
For the Month ending March 31, 2016
Indirect factory wages
$86,000
Depreciation of plant and e

Therefore, total factory overhead is $104,300.

1.6)

Computation of cost of goods sold budget:

$19.480
G Co
Cost of Goods sold Budget
For the month ending March 31, 2016
Finished goods inventory, March 1, 2016
Work in

1.7)

Computation of selling and administrative expense budget for March:

G Co
Selling and Administrative expenses Budget
For the month ending March 31, 2016
Selling expense:
Sales salaries expense

Therefore, total operating expense is $325,800.

1.8)

Computation of Budgeted income statement for the month ending March 31, 2016:

G Co
Budgeted Income statement
For the month ending March 31, 2016
Particulars Amount Amount
Revenue from sales $1.088.000

Therefore, net income is $160,230.

Workings:

Calculation of total revenue from sales:

Revenuefromsales=(Revenuefrombattinghelmet+Revenuefromfootballhelmet)=$48,000+$1,040,000=$1,088,000\begin{array}{c}\\{\rm{Revenue from sales = }}\left( {{\rm{Revenue from batting helmet + Revenue from football helmet}}} \right)\\\\ = \$ 48,000 + \$ 1,040,000\\\\ = \$ 1,088,000\\\end{array}

Therefore, revenue from sales is $1,088,000.

Calculation of revenue from batting helmet and football helmet:

Salesrevenue=Unitssold×SellingpriceperunitBattinghelmet=1,200×$40=$48,000Footballhelmet=6,500×$160=1,040,000\begin{array}{c}\\{\rm{Sales revenue = Units sold}} \times {\rm{Selling price per unit}}\\\\{\rm{Batting helmet = 1,200}} \times {\rm{\$ 40 }}\\\\{\rm{ = \$ 48,000}}\\\\{\rm{Football helmet = 6,500}} \times {\rm{\$ 160}}\\\\{\rm{ = 1,040,000}}\\\end{array}

Therefore revenue from batting helmet and football helmet is $48,000 and $1,040,000.

Calculation of income tax:

Incometax=Incomebeforeincometax×30%=$228,900×30%=$68,670\begin{array}{c}\\{\rm{Income tax = Income before income tax}} \times 30\% \\\\ = \$ 228,900 \times 30\% \\\\ = \$ 68,670\\\end{array}

Therefore, income tax is $68,670.

Ans: Part 1.2

G Co
Production Budget
For the Month ending March 31, 2016
Particulars
Units
Batting helmet Football helmet
Expected units

Part 1.3

G Co
Material Purchase Budget
For the Month ending March 31, 2016
Particulars
Plastic Foam Lining Total
Units required for

Part 1.4

G Co
Direct Labor cost Budget
For the Month ending March 31, 2016
Molding Assembly Total
Hours required for production
Batt

Part 1.5

G Co
Factory overhead Budget
For the Month ending March 31, 2016
Indirect factory wages
$86,000
Depreciation of plant and e

Part 1.6

G Co
Cost of Goods sold Budget
For the month ending March 31, 2016
19.480
15,300
Finished goods inventory, March 1, 2016
Wo

Part 1.7

$9,000
G Co
Selling and Administrative expenses Budget
For the month ending March 31, 2016
Selling expense:
Sales salaries

Part 1.8

G CO
Budgeted Income statement
For the month ending March 31, 2016
Particulars
Amount Amount
Revenue from sales $1,088,000